OpenStudy (anonymous):

A solution contains 12.5% NaCl by mass. What mass of solution is required to obtain 1.0 x 10^23 Na atoms?

OpenStudy (anonymous):

can you make a conversion factor out of the percent by mass?

OpenStudy (anonymous):

no I think that its has something to do with problem solving like 1.0 x 10^23 with density= volume/mass. Or another formula im not to sure i am stumped at this problem sorry.

OpenStudy (anonymous):

alright I'm going to show you how to do this factor label style because once you get this down these things are no brainers. The first thing we know that is a conversion factor is 12.5% nacl by mass. that means that for every 100g of solution we get 12.5g of nacl. this is a conversion factor 12.5g NaCl=100g solution

OpenStudy (anonymous):

This allows us to go from NaCl to solution or solution to NaCl depending on how we postion it in the factor label

OpenStudy (anonymous):

You don't really need this part however it is good to see where it comes from so you will now need a balanced ionic equation. \[Na^+ + Cl^- \rightarrow NaCl\] What can you tell me about the stoichiometry of this equation?

OpenStudy (anonymous):

what I can tell is that an electron left Na and added itself to Cl making Cl-, while making Na+ which makes the compound NaCl

OpenStudy (anonymous):

Ok that is good and is correct but not what we are looking for at this point in time. Have you ever heard of the term stoichiometry?

OpenStudy (anonymous):

No sorry I haven't.

OpenStudy (anonymous):

If not it really isn't much more than using the coefficient in front of the substances in your balanced chemical equation. In this case everything is 1 to 1. This is going to be helpful when we set up or factor label.

OpenStudy (anonymous):

Ok can you tell me the molar mass of NaCl

OpenStudy (anonymous):

is it 58.4425 g/mol

OpenStudy (anonymous):

\[1.0x10^{23} Na^+ atoms (\frac{1 mole Na^+ }{6.022x10^{23} Na+ atoms})(\frac{1 mole NaCl}{1 mole Na^+})(\frac{58.44 g NaCl}{1mole NaCl})(\frac{100g Solution}{12.5g NaCl})=?\]

OpenStudy (anonymous):

have you ever seen anything like that?

OpenStudy (anonymous):

and if I did my math right you end up needing 77.64g of solution, I didn't check your molar mass but used it in a conversion factor so if that is off it will change the answer.

OpenStudy (anonymous):

I have seen something like that, but I think I do it more different than that. I know that the mole is 6.022x10^22 was somewhere inside the equation but I always get mixed up when doing these types of equations.

OpenStudy (anonymous):

Thank you for showing me how to do the problem.

OpenStudy (anonymous):

No problem I hope it helped one more trick to make these go easier is to make an arrow plan. It allows you to decide the units you are going to next using my method I will give you an example using this problem. But basically determine the units you want to start with and then end up with. Then determine the conversion factors that help you get to the goal.

OpenStudy (anonymous):

ok. thank you zbay.

OpenStudy (anonymous):

\[Na^+ atom \rightarrow mol Na \rightarrow mol NaCl \rightarrow g NaCl \rightarrow g solution\] now that you have a visual plan it may be easier to determine what conversion factors to use. Good luck and hope this helps!

OpenStudy (anonymous):

Thank you for helping me.